OpenStudy (anonymous):

Can someone please help me graph this function? f(x)=sqrt(x+3)-2

jhonyy9 (jhonyy9):

for example give for x the value of 1 so than you will get for y =sqrt(1+3) -2 and for x=6 than y=sqrt(6+3) -2 so calculi the values of y and just graph these points hope that this will help you

OpenStudy (anonymous):



OpenStudy (anonymous):

i)draw sq. root x graph

OpenStudy (anonymous):

ii) by shifting x- axis by 3

OpenStudy (anonymous):

in left side then shift by 2 in downward direction

OpenStudy (anonymous):

hope this helps you

OpenStudy (anonymous):

Thanks so much!